OpenStudy (cutestgirlaround):

PLEASE HELP!!! What value of a makes the equation true? 1.7a+0.3a=4/5 A. a=1 3/5 B. a=2 4/5 C. a=−1 1/5 D. a=2/5

OpenStudy (qwertty123):

What do you think?

OpenStudy (cutestgirlaround):

C or D

OpenStudy (sapphiremoon):

Use your calculator to substitute each value into the equation.

OpenStudy (sapphiremoon):

Or, you can try adding like terms and solving for a.

OpenStudy (cutestgirlaround):

I still think C or D

OpenStudy (cutestgirlaround):

I know A and B are to high but I'm leaning more towards D

563blackghost (563blackghost):

Combine the like terms.... \(\huge{(1.7a+0.3a) \rightarrow 2a}\) In which we have this equation.... \(\huge{2a=\frac{4}{5}}\) We would then divide by 2 to get the variable a to itself.... \(\huge{\frac{2a}{2}=\frac{4}{5} \div 2 \rightarrow a = \frac{4 \div 2}{5}}\)

563blackghost (563blackghost):

So yea you are correct about it being D more :D

OpenStudy (cutestgirlaround):

thanks :-D

563blackghost (563blackghost):

np :)